OpenStudy (malikasif):

Verify by defination limit x approches infinty (x^2+4x+8)/(2x^3-3x^2+2) =1/2

OpenStudy (irishboy123):

\[\dfrac{(x^2+4x+8)}{(2x^3-3x^2+2)} \] you sure it's not \[\dfrac{(x^\color{red}{3}+4x+8)}{(2x^3-3x^2+2)} \]

OpenStudy (malikasif):

Yeah i am sure

OpenStudy (abmon98):

Divide by the largest power on the denominator lim x-->infinity(1/x+4/x^2+8/2-3/x+2/x^3)=0 you wont get 1/2 unless its x^3 in the numerator.

OpenStudy (irishboy123):

yes, Ab !!

OpenStudy (welshfella):

thats correct

OpenStudy (crabbyoldgamer):

The limit as x approaches infinity of (x^2+4x+8)/(2x^3-3x^2+2) = 0